I need assist on this:
2(12-8X)=10X
4(X-2)=5X+2

Answer by sachi(548)

I need assist on this:
2(12-8X)=10X
4(X-2)=5x+2
ans:-
2(12-8X)=10X
or 24-16x=10x
or 10x+16x=24
or 26x=24
or x=24/26=12/13 ans
4(X-2)=5x+2
or 4x-8=5x-2
or 5x-4x=-8+2
or x=-6 ans